3. a. Identify the limiting reagent when 15.0 g of H2 reacts with 48.0 g of O2 to produce water. b. What is the mass of water produced?
Answer #1

3.

a)

Balanced reaction is

H2 + O2 H2O

So, 1 mole Hydrogen reacts with mole oxygen .

Given , Mass of Hydrogen = 15 g,

so moles of Hydrogen =(15/2) = 7.5

Hence, moles of oxygen required for complete reaction

= (7.5/2) = 3.75

But, Mass of oxygen taken = 48 g,

hence, moles of oxygen = (48/32) = 1.5

As moles of oxygen taken is less than required, so limiting reagent is oxygen.

B)

Moles of water formed = 2* moles of oxygen reacted.

= 2*1.5 = 3 .

Mass of 3 mole water = 3*18= 54 g.